OpenStudy (anonymous):

Which point is a solution of x + 2y ≤ 4?

OpenStudy (campbell_st):

well there will be numerous points that satisfy the inequality. You should substitute each ordered pair from your answer list the correct solution will be the ordered pair that is true. e.g. (-10, -10) is \[-10 + 2 \times (-10) \le 4\] this is true because \[-30 \le 4\] so that is the process you need to work through